OpenStudy (anonymous):

I need help with this question 9x - 2y =-54 7x + 47 = y substitution method

OpenStudy (anonymous):

substitute y=7x+47 in equation one first

OpenStudy (anonymous):

The second line to be reduced to x+7=y, so the first line is... 9x-2(x+7)=-54 9x-2x-14=-54 add 14 to each side 9x-2x=40 7x=40

OpenStudy (anonymous):

x=40/7

OpenStudy (anonymous):

small mistake misspacgirl in line 2,-2(x+47):)

OpenStudy (anonymous):

OHHHHHH... oopsss! Im sryyy! :O

OpenStudy (anonymous):

x=-8 y=-9

OpenStudy (anonymous):

one small mistake can make a huge difference!!!